The length of a rectangle is 1 cm longer than the width. If the diagonal of the rectangle is 4cm, what are the dimensions (the length and the width) of the rectangle.
Please show work.

Answer by checkley75(3666)

THE DIAGONAL FORMS A TRIANGLE WITH SIDES X, X+1 & 4
USING THE PATHOGOREAN FORMULA
A^2+B^2=C^2
X^2+(X+1)^2=4^2
X^2+X^2+2X+1=16
2X^2+2X+1-16=0
2X^2+2X-15=0
2X^2+2X-15=0
USING THE QUADRATIC EQUATION WE GET
X=(-2+-SQRT[2^2-4*2*-15])/2*2
X=(-2+-SQRT4+120])/4
X=(-2+-SQRT124)/4
X=(-2+-11.13)/4
X=(-2+11.13)/4
X=9.13/4
X=2.28 ANSWER.
X=(-2-11.13)/4
X=-13.13/4
X=-3.28 ANSWER.